OpenStudy (anonymous):

Solutions for -3x+2y=8

OpenStudy (zale101):

find the x-intercept, set y to 0: -3x + 2(0) = 8 -3x + 0 = 8 -3x = 8 x = ? Now try to find the y-intercept, set x to 0

OpenStudy (anonymous):

the y intercept is 4 is it not?

OpenStudy (zale101):

To find the y-intercept, set x to 0: -3(0) + 2y = 8 0 + 2y = 8 2y = 8 y = 8/2 = 4 correct!

OpenStudy (anonymous):

and the x intercept is -8/3

OpenStudy (zale101):

x=-8/3 y=4

OpenStudy (zale101):

correct correct correct :D

OpenStudy (anonymous):

thank you

OpenStudy (zale101):

no problem :)